OpenStudy (anonymous):

Sam is observing the velocity of a car at different times. After two hours, the velocity of the car is 50 km/h. After six hours, the velocity of the car is 54 km/h. Part A: Write an equation in two variables in the standard form that can be used to describe the velocity of the car at different times. Show your work and define the variables used. Part B: How can you graph the equation obtained in Part A for the first seven hours?

OpenStudy (anonymous):

use slope formula (54-50)/(6-2) so m=1

OpenStudy (anonymous):

then use point slope form : y-50=x-2 so the answer should be x-y=-48

OpenStudy (anonymous):

Is that A and B?

OpenStudy (anonymous):

thats A only

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

Can you help with B too please?

OpenStudy (anonymous):

for B Plug in 0 for t and then 6 for t...you get (0,50) and (6,62) Connect those 2 points and you have your graph for the first 6 hours

OpenStudy (anonymous):

Thanks bro!!!

OpenStudy (anonymous):

no prob